### v1

# Spectral Properties of Irradiated Circumbinary Disks around Binary Black Holes Governed by Hydrogen Opacities Dependent on Temperature and Density

Abstract
We study the thermal and spectral properties of irradiated circumbinary disks (CBDs) around binary black holes (BBHs), using analytic, hydrogen-based opacity models that capture key dependencies on temperature, density, and ionization. We solve the vertically hydrostatic energy balance equations with Rosseland mean opacities from free-free absorption, bound-free absorption, and electron scattering processes, combined with ionization fractions derived by the Saha equation. Four opacity models are considered, including a reference model with no physical opacity, constructed by Lee et al. (2024), and three physically motivated alternatives. The midplane temperature profiles show significant variation across models, while the surface temperature remains largely unchanged in regions dominated by viscous heating. Opacity effects become pronounced in the outer disk, where irradiation reprocessing shapes the IR-optical continuum. Inclusion of bound-free opacity introduces a noticeable flattening and a mid-frequency peak in the spectral energy distribution. We also compute spectra of a triple disk system including the CBD and two accreting minidisks. The high-frequency peak arises from the hot minidisks, while the low-frequency excess originates from irradiated outer CBD layers. By comparing model spectra with detection limits of Subaru, JWST, and Swift, we find that systems within $\sim 10$ Mpc may show detectable IR excess from the CBD. Our results highlight the need for accurate opacity modeling to interpret electromagnetic signatures of black hole mergers and support future integration of opacity tables with metallicity.
accretion, accretion discs – black hole physics – gravitational waves – hydrodynamics – binaries: general
1 Introduction
Stellar-mass binary black holes (BBHs), composed of two black holes with masses in the range ${3\,M_{\odot}}\leq M\leq{100\,M_{\odot}}$, can form either through the isolated binary evolution of massive stars (Marchant et al., 2016) or through dynamical interactions in dense stellar systems such as globular clusters (Morscher et al., 2015). Since the first direct detection of a black hole merger by the LIGO Observatory (event GW150914) (Abbott et al., 2016), a growing number of such events have been reported (Abbott et al., 2023; Nitz et al., 2023), including those suggesting the existence of intermediate-mass black holes (IMBHs). On a larger scale, supermassive black hole (SMBH) binaries are thought to reside at the centers of active galactic nuclei (AGNs) and may play a key role in galaxy evolution as well as structure formation in the early universe. At all these mass scales, the presence of circumbinary disks (CBDs) around BBHs provides an important way to probe binary evolution via electromagnetic (EM) signals complementary to gravitational waves (GWs).
The thermal and spectral properties of a CBD depend critically on its internal heating and radiative cooling processes. In particular, CBDs are expected to be heated by viscous dissipation and, in some cases, by irradiation from an accreting minidisk around each black hole. A recent study by Lee et al. (2024) investigated the effect of irradiation from such minidisks in shaping the CBD temperature profile and spectrum, showing that reprocessed emission can lead to a characteristic infrared (IR) peak in the spectral energy distribution (SED). However, their analysis assumed the disk surface temperature equals to the disk midplane temperature, resulting in a fixed optical depth $\tau\lesssim 1$, thereby neglecting the strong temperature and ionization dependence of opacity in cooler, denser regions of the CBD.
In reality, the opacity of a hydrogen-rich disk varies significantly with local temperature $T$, density $\rho$, and ionization degree $x_{e}(T)$ (Rybicki & Lightman, 1979). In particular, for the midplane temperature of an accretion disk in the range $10^{3}~{}{\rm K}\lesssim T\lesssim 10^{4}~{}{\rm K}$, the dominant sources of opacity are known to be bound-free ($\kappa_{\rm bf}$) and free-free ($\kappa_{\rm ff}$) absorption, while electron scattering ($\kappa_{\rm es}$) typically dominates for $T\gtrsim 10^{4}\,{\rm K}$ (e.g., Shakura & Sunyaev 1973; Kato et al. 2008). These absorption opacities are highly sensitive to $T$ and $x_{e}$, the latter of which is determined by the Saha equation (Saha, 1920).
In this paper, we examine the impact of different opacity sources on the thermal structure and the spectrum of an irradiated CBD. We focus on three opacity prescriptions: $\kappa_{\rm es}$, $\sqrt{\kappa_{\rm es}\kappa_{\rm ff}}$, and $\sqrt{\kappa_{\rm a}(\kappa_{\rm a}+\kappa_{\rm es})}$, where $\kappa_{\rm a}=\kappa_{\rm ff}+\kappa_{\rm bf}$, and assess how they affect the radial profiles of the CBD aspect ratio $H/r$, the CBD midplane temperature $T_{\rm c}$, the surface temperature $T_{\rm s}$, and the spectrum emanating from the CBD surface.
For this purpose, we use Rosseland-mean opacities computed from analytic Kramers’ formulas for hydrogen bound-free and free-free absorption processes, incorporating the ionization degree calculated from the Saha equation. We solve the energy balance equation,
$Q_{\rm vis}+Q_{\rm irr}=Q_{\rm rad},$ (1)
where $Q_{\rm vis}$ is the viscous heating rate, $Q_{\rm irr}$ is the irradiation heating rate, and $Q_{\rm rad}$ is the radiative cooling rate depending on $\kappa(\rho,T_{\rm c},x_{e}(T_{\rm c}))$. The resulting models allow us to quantify how microphysical opacity prescriptions shape the thermal state and spectral signature of the CBD with and without irradiation.
Our study aims to establish a physically consistent baseline for understanding irradiated CBDs in a low temperature regime and to facilitate future comparisons with models incorporating metallicity, dust, or full opacity tables. These results may also aid in the interpretation of EM counterparts to black hole mergers or merging BHs detected via GWs.
In Section 2, we present the basic framework for modeling a one-dimensional (1D), geometrically thin, optically thick, irradiated CBD with the detailed description of the three opacity prescriptions. In Section 3, we numerically solve the energy equation of the CBD. Among the three opacity models, the model involving combined free-free absorption, bound-free absorption, and electron scattering processes requires numerical integration due to their nonlinear dependence on the density, temperature, and ionization degree. In Section 4, we present the resulting radial profiles of disk temperature, the emergent spectra from the CBD, and the combined spectra from the CBD and minidisks, referred to as the triple disk spectra. Sections 5 and 6 are devoted to discussion and conclusions, respectively.
Figure 1: Schematic illustration of an irradiated circumbinary disk (CBD) surrounding two minidisks orbiting the primary and secondary black holes. The minidisks, referred to as the circum-primary disk (CPD) and the circum-secondary disk (CSD), are gravitationally bound to the primary (BH1) and secondary (BH2) black holes, respectively. The yellow ochre, blue, and red shaded regions represent the CBD, CPD, and CSD, respectively. The two black circles at the center denote the black holes, with their common center of mass (CM) marked by a plus sign. The blue and red wavy arrows indicate the trajectories of irradiating photons emitted from the minidisks. The inner edge of the CBD is located at a radius $r_{\rm in}=C_{\rm gap}a$, where $C_{\rm gap}$ is a parameter characterizing the size of the gap between the CBD and the binary, and $a$ is the binary’s semi-major axis.
2 Model
We consider a 1D, axisymmetric, and steady-state CBD surrounding a BBH system. The CBD is assumed to be optically thick and geometrically thin, and to evolve under viscous dissipation in the presence of external irradiation from the minidisks.
Our goal is to investigate how the thermal structure of the CBD and its emergent spectrum respond to different physical forms of the Rosseland mean opacity, particularly in the midplane temperature range $10^{3}\,{\rm K}\lesssim T_{\rm c}\lesssim 10^{5}\,{\rm K}$, where temperature- and ionization-dependent hydrogen absorption dominates. To isolate this dependence, we assume a pure hydrogen composition.
The total mass of the binary is defined as $M=M_{1}+M_{2}$, where $M_{1}$ and $M_{2}$ are the masses of the primary and secondary black holes, respectively. The binary follows a circular orbit with a semi-major axis $a=a_{0}\,r_{\rm S}$, where $a_{0}$ is a dimensionless scaling factor and the Schwarzschild radius, $r_{S}$, is defined as
$r_{\rm S}$ $=\frac{2GM}{c^{2}}$ (2)
$\sim 3.0\times 10^{7}\,{\rm cm}\left(\frac{M}{100\,M_{\odot}}\right),$
where $G$ is the gravitational constant and $c$ is the speed of light. As illustrated in Figure 1, the CBD inner-edge radius is given by
$r_{\rm in}$ $=C_{\rm gap}a$ (3)
$=6.0\times 10^{10}\,{\rm cm}\left(\frac{C_{\rm gap}}{2}\right)\left(\frac{a_{0}}{1000}\right)\left(\frac{M}{100\,M_{\odot}}\right)$
with $1.6\lesssim C_{\rm gap}\lesssim 4$ (Artymowicz & Lubow, 1994; Dittmann et al., 2023). We adopt $C_{\rm gap}=2$ as a fiducial value throughout this paper.
It is noted that irradiation originating near the inner edge of each minidisk is unlikely to affect the spectrum of the outer region, since the radius at which irradiation heating dominates over viscous heating is well beyond the disk outer edge of each minidisk. In addition, photons emitted from the inner edge of the CBD are primarily in the optical band and have significantly lower energies than those from the inner edges of the minidisks. This difference arises because the temperatures of the minidisks are higher than the CBD temperature, following the standard disk relation $T\propto r^{-3/4}$ (Pringle, 1981). Consequently, the contribution of irradiation from the CBD’s inner edge to the overall SED is negligible. We therefore omit the irradiation heating rate from the CBD inner edge in the energy equation.
The radiative cooling rate from both sides of the disk surface is given by
$Q_{\rm rad}=2\sigma T_{\rm s}^{4},$ (4)
where $T_{\rm s}$ is the surface temperature and $\sigma$ is the Stefan-Boltzmann constant. Assuming local radiative equilibrium in the vertical direction of the CBD, the radiative flux can also be expressed as
$F_{\rm rad}=\frac{32}{3}\frac{\sigma T_{\rm c}^{4}}{\tau},$ (5)
where $T_{\rm c}$ is the midplane temperature and
$\tau=\frac{1}{2}\kappa(\rho,T_{\rm c},x_{e})\Sigma$ (6)
is the optical depth, with $\Sigma$ being the surface density of the disk. In our formulation, the opacity $\kappa(\rho,T_{\rm c},x_{e})$ consists of a sum of contributions from electron scattering, free-free absorption, and bound-free absorption (see Section 3 for details). Equating equation (4) with equation (5) yields the relation between the surface and midplane temperatures:
$T_{\rm s}=\frac{2}{3^{1/4}}\frac{1}{\tau^{1/4}}T_{\rm c}.$ (7)
According to the standard disk theory, the heating rate due to the viscous heating of the disk is given by
$Q_{\rm vis}=\frac{3GM\dot{M}}{4\pi r^{3}}$ (8)
for $r\gg{r_{\rm ISCO}}$, where $r_{\rm ISCO}=3\,r_{\rm S}$ is the radius at the innermost stable circular orbit (ISCO) around a Schwarzschild black hole of mass $M$. We parameterize $\dot{M}$ in terms of the Eddington accretion rate as
$\dot{M}=\dot{m}\frac{L_{\rm Edd}}{c^{2}},$ (9)
where $\dot{m}$ denotes the dimensionless accretion rate, which is set to unity throughout this paper, and
$L_{\rm Edd}=\frac{4\pi GMc}{\kappa_{\rm es}}$ (10)
is the electron-scattering based Eddington luminosity. The electron-scattering opacity is set to be $\kappa_{\rm es}=0.2(1+X)\,{\rm cm^{2}\,g^{-1}}$, where $X$ is the hydrogen mass fraction. In the following we assume $X=1$ and thus $\kappa_{\rm es}=0.4\,{\rm cm^{2}\,g^{-1}}$.
The disk viscosity is given by (Shakura & Sunyaev, 1973)
$\nu=\frac{2}{3}\alpha_{\rm SS}c_{\rm s}H$ (11)
with the viscosity parameter $\alpha_{\rm SS}$, where $c_{\rm s}$ is the sound speed. Assuming the CBD is in hydrostatic equilibrium for the direction perpendicular to the disk plane, the sound speed $c_{\rm s}$ is given by
$c_{\rm s}=\Omega{H},$ (12)
where $\Omega=\sqrt{GM/r^{3}}$ is the Keplerian angular frequency of the CBD.
The ratio between the CBD viscous timescale $\tau_{\rm vis}(r)=r^{2}/\nu$ at the CBD inner edge and the binary orbital period $P_{\rm orb}$ is given by using equations (11) and (12) as
$\frac{\tau_{\rm vis}\left(r_{\rm in}\right)}{P_{\rm orb}}$ $\sim$ $2.0\times 10^{4}\left(\frac{C_{\rm gap}}{2}\right)^{3}\left(\frac{a_{0}}{1000}\right)^{3/2}\left(\frac{\alpha_{\rm SS}}{0.1}\right)^{3}\left(\frac{H/r}{0.01}\right)^{2},$ (13)
where $P_{\rm orb}=2\pi/\Omega\sim 2.8\times 10^{2}\,{\rm s}\,\left(a_{0}/1000\right)^{3/2}\left(M/100M_{\odot}\right)^{-1/2}$.
Since equation (13) suggests that $\tau_{\rm vis}\gg P_{\rm orb}$, we assume that the binary system with triple disk composed of CBD, CPD, and CSD is in a quasi-steady state. This allows to impose on the following relation on the accretion rates between CBD, CPD, and CSD:
$\dot{M}=\dot{M}_{1}+\dot{M}_{2},$ (14)
where $\dot{M}_{1}$ and $\dot{M}_{2}$ are the CPD and CSD accretion rates, respectively. Also, we assume that the mass accretion rate ratio is equal to the binary mass ratio $q$, i.e. $\dot{M}_{2}/\dot{M}_{1}=q$. These relations provide
$\dot{M}_{1}=\frac{1}{1+q}\dot{M},\,\,\dot{M}_{2}=\frac{q}{1+q}\dot{M}.$ (15)
The bolometric luminosities of the primary and secondary black holes are given by
$L_{1}=\frac{1}{6}\dot{M}_{1}c^{2},\,\,L_{2}=\frac{1}{6}\dot{M}_{2}c^{2}.$ (16)
The equation of state for an ideal gas gives the sound speed as
$c_{\rm s}=\sqrt{\frac{R_{\rm g}}{\mu}T_{\rm c}},$ (17)
where $R_{\rm g}$ is the gas constant and $\mu$ is the mean molecular weight. For hydrogen gas, $\mu$ is given by
$\mu=\frac{1}{1+x_{e}(T_{\rm c})},$ (18)
where $x_{e}(T_{\rm c})$ denotes the ionization degree, where $x_{e}=0$ and $x_{e}=1$ correspond to neutral hydrogen atoms and fully ionized hydrogen gas, respectively. The ionization fraction is determined by the Saha equation (Saha, 1920):
$\frac{x_{e}^{2}}{1-x_{e}}=2\left(\frac{2\pi m_{e}kT_{\rm c}}{h^{2}}\right)^{3/2}\frac{1}{n_{\rm H}}\frac{g_{\rm i}}{g_{\rm n}}\exp\left(-\frac{\chi}{kT_{\rm c}}\right),$ (19)
where $m_{e}$ is the electron mass, $h$ is the Planck constant, $n_{\rm H}$ is the number density of hydrogen atoms, $k$ is the Boltzmann constant, $g_{\rm i}$ and $g_{\rm n}$ are the statistical weights of the ionized and neutral states, respectively, and $\chi$ is the ionization energy. Throughout this paper, we adopt $g_{\rm i}=1$, $g_{\rm n}=1$, and $\chi=13.6\,{\rm eV}$.
Equating equation (12) with equation (17) gives the CBD midplane temperature:
$T_{\rm c}=\frac{1}{R_{\rm g}}\frac{1}{1+x_{e}}\left(\frac{H}{r}\right)^{2}\frac{GM}{r}.$ (20)
2.1 Disk Opacity
In partially ionized hydrogen gas, the dominant sources of opacity are free-free absorption ($\kappa_{\rm ff}$), bound-free absorption ($\kappa_{\rm bf}$), and electron scattering ($\kappa_{\rm es}$). The free-free opacity arises from the bremsstrahlung absorption by electrons interacting with protons, and is given approximately by (e.g., Rybicki & Lightman 1979)
$\kappa_{\rm ff}=\kappa_{\rm ff,0}x_{e}^{2}\rho\,T_{\rm c}^{-7/2}\,{\rm cm^{2}g^{-1}},$ (21)
where $\kappa_{\rm ff,0}=1.1\times 10^{23}\,{\rm cm^{5}g^{-2}K^{7/2}}$ and $\rho$ is the mass density in g cm-3. The bound-free opacity, caused by photoionization of neutral hydrogen atoms, is typically written as
$\kappa_{\rm bf}=\kappa_{\rm bf,0}(1-x_{e})\rho\,T_{\rm c}^{-7/2}\,{\rm cm^{2}g^{-1}},$ (22)
where $\kappa_{\rm bf,0}=4.3\times 10^{25}\,{\rm cm^{5}g^{-2}K^{7/2}}$ (e.g., Rybicki & Lightman 1979). While $\kappa_{\rm bf}$ is most effective in the temperature range $T_{\rm c}\sim 5000$-$10^{4}$ K, it can still be applied down to $T_{\rm c}\sim 1000$-$5000$ K in hydrogen-rich accretion disks.
However, in realistic astrophysical conditions, especially at $T\lesssim 5000$ K, the true opacity deviates significantly from the hydrogen $\kappa_{\rm bf}$ because of contributions from metals and dust grains (e.g., Semenov et al. 2003; Ferguson et al. 2005). These include the line and continuum absorption by molecules and solid particles, which can increase the opacity by orders of magnitude. Despite this, using $\kappa_{\rm bf}$ derived from pure hydrogen models remains useful for highlighting the role of hydrogen ionization and providing an upper bound on the opacity in irradiated disks. This approach is particularly appropriate for idealized accretion disks composed of hydrogen, enabling us to isolate fundamental physical effects without introducing chemical complexity.
The effective opacities used in our model are given by (Rybicki & Lightman, 1979)
$\kappa_{1}$ $=$ $\kappa_{\rm es},$ (23)
$\kappa_{2}$ $=$ $\sqrt{\kappa_{\rm es}\kappa_{\rm ff}},$ (24)
$\kappa_{3}$ $=$ $\sqrt{\kappa_{\rm a}\left(\kappa_{\rm a}+\kappa_{\rm es}\right)},$ (25)
where the absorption opacity $\kappa_{\rm a}$ is given by
$\kappa_{\rm a}=\kappa_{\rm ff}+\kappa_{\rm bf}.$ (26)
To explore how different opacity prescriptions affect the thermal structure and emergent spectrum of the CBD, we consider four representative models. Model 0 reproduces the fiducial case of Lee et al. (2024), in which no physical source of opacity is included by assuming that $T_{\rm s}=T_{\rm c}$. This model serves as a baseline for comparison with the three physically motivated opacity models. Model 1 includes only electron scattering opacity, representing the fully ionized regime where scattering dominates. Model 2 assumes a hybrid regime where electron scattering is stronger than but not entirely dominant over free-free absorption. Model 3 employs a more integrated treatment by accounting for both the two absorption processes and electron scattering. Table 1 summarizes the adopted opacity prescriptions, associated optical depth expressions, and boundary conditions for each model.
3 Method
Solving the energy equation yields the disk aspect ratio $H/r$, which can then be substituted into the hydrostatic equation to finally obtain the radial distribution of the CBD temperature. In the following, we describe the method to solve the energy equation numerically. First, we introduce the following dimensionless variables:
$\xi$$\equiv$ $\frac{r}{r_{\rm in}}\hskip 7.11317pt{\rm and}\hskip 7.11317ptY\equiv\frac{H}{r}.$ (27)
Equation (20) is rewritten with equation (27) as
$T_{\rm c}(1+x_{e}(T_{c}))=T_{0}\frac{Y^{2}}{\xi},$ (28)
where we define the dynamical temperature at $r_{\rm in}$ as
$T_{0}$ $=\frac{1}{R_{g}}\frac{GM}{r_{\rm in}}\sim 2.7\times 10^{9}\,{\rm K}\left(\frac{C_{\rm gap}}{2}\right)^{-1}\left(\frac{a_{0}}{1000}\right)^{-1}.$
3.1 Optical depth prescriptions
Considering that the CBD is modeled by the standard disk model, the surface density is given by
$\Sigma$ $=$ $\frac{\dot{M}}{3\pi\nu}=\sqrt{2}\frac{\dot{m}}{\alpha_{\rm SS}}\frac{1}{\sqrt{C_{\rm gap}a_{0}}}\frac{1}{\kappa_{\rm es}}\frac{1}{Y^{2}\xi^{1/2}},$ (29)
where we used equations (3), (9), (11), (12) and (27) for the derivation. The mass density is then given by
$\rho=\frac{1}{2}\frac{\Sigma}{H}$ $=$ $\frac{1}{2}\frac{r}{H}\frac{\Sigma}{r_{\rm S}}\frac{r_{\rm S}}{r_{\rm in}}\frac{r_{\rm in}}{r}=\frac{1}{2}\frac{1}{Y\xi}\frac{\Sigma}{r_{\rm S}}\frac{1}{C_{\rm gap}a_{0}}=\frac{1}{\sqrt{2}}\frac{\dot{m}}{\alpha_{\rm SS}}\frac{1}{(C_{\rm gap}a_{0})^{3/2}}\frac{1}{r_{\rm S}\kappa_{\rm es}}\frac{1}{Y^{3}\xi^{3/2}}.$ (30)
Substituting equations (28) and (30) into equations (21) and (22) yields the free-free and bound-free absorption opacities in the form
$\kappa_{\rm ff}$ $=$ $\kappa_{\rm ff,0}\,\rho\,T_{\rm c}^{-7/2}$
$=$ $\frac{1}{\sqrt{2}}\frac{\dot{m}}{\alpha_{\rm SS}}\frac{1}{(C_{\rm gap}a_{0})^{3/2}}\left[\frac{\kappa_{\rm ff,0}}{\kappa_{\rm es}}\frac{x_{e}^{2}}{r_{\rm S}}\left(\frac{1+x_{e}}{T_{0}}\right)^{7/2}\right]\frac{\xi^{2}}{Y^{10}},$
$\kappa_{\rm bf}$ $=$ $\kappa_{\rm bf,0}\,\rho\,T_{\rm c}^{-7/2}$
$=$ $\frac{1}{\sqrt{2}}\frac{\dot{m}}{\alpha_{\rm SS}}\frac{1}{(C_{\rm gap}a_{0})^{3/2}}\left[\frac{\kappa_{\rm bf,0}}{\kappa_{\rm es}}\frac{1-x_{e}}{r_{\rm S}}\left(\frac{1+x_{e}}{T_{0}}\right)^{7/2}\right]\frac{\xi^{2}}{Y^{10}}.$
Equation (26) can be rewritten as
$\kappa_{\rm a}$ $=$ $\frac{1}{\sqrt{2}}\frac{\dot{m}}{\alpha_{\rm SS}}\frac{1}{(C_{\rm gap}a_{0})^{3/2}}\frac{\kappa_{\rm ff,0}}{\kappa_{\rm es}}\frac{1}{r_{\rm S}}x_{e}^{2}\left(\frac{1+x_{e}}{T_{0}}\right)^{7/2}\left[1+\frac{1-x_{e}}{x_{e}^{2}}\frac{\kappa_{\rm bf,0}}{\kappa_{\rm ff,0}}\right]\frac{\xi^{2}}{Y^{10}}.$ (33)
Equations (24) and (25) are also rewritten as
$\kappa_{2}$ $=$ $\frac{1}{2^{1/4}}\left(\frac{\dot{m}}{\alpha_{\rm SS}}\right)^{1/2}\frac{1}{(C_{\rm gap}a_{0})^{3/4}}\left(\frac{\kappa_{\rm ff,0}}{r_{\rm S}}\right)^{1/2}x_{e}\left(\frac{1+x_{e}}{T_{0}}\right)^{7/4}\frac{\xi}{Y^{5}},$ (34)
$\kappa_{3}$ $=$ $\frac{1}{\sqrt{2}}\frac{\dot{m}}{\alpha_{\rm SS}}\frac{1}{(C_{\rm gap}a_{0})^{3/2}}\frac{\kappa_{\rm ff,0}}{\kappa_{\rm es}}\frac{1}{r_{\rm S}}\left(\frac{1+x_{e}}{T_{0}}\right)^{7/2}\left[x_{e}^{2}+(1-x_{e})\frac{\kappa_{\rm bf,0}}{\kappa_{\rm ff,0}}\right]\frac{\xi^{2}}{Y^{10}}$ (35)
$\times$ $\Biggr{[}1+\sqrt{2}\frac{\alpha_{\rm SS}}{\dot{m}}(C_{\rm gap}a_{0})^{3/2}r_{\rm S}\frac{\kappa_{\rm es}^{2}}{\kappa_{\rm ff,0}}\left(\frac{T_{0}}{1+x_{e}}\right)^{7/2}\left[\frac{1}{x_{e}^{2}+(1-x_{e})(\kappa_{\rm bf,0}/\kappa_{\rm ff,0})}\right]\frac{Y^{10}}{\xi^{2}}\Biggr{]}^{1/2}.$
Then, the optical depth for each opacity is given by
$\tau_{1}$ $=$ $\frac{1}{2}\kappa_{1}\Sigma=\tau_{10}\frac{1}{Y^{2}\xi^{1/2}},$ (36)
$\tau_{2}$ $=$ $\frac{1}{2}\kappa_{2}\Sigma=\tau_{20}\frac{\xi^{1/2}}{Y^{7}},$ (37)
$\tau_{3}$ $=$ $\frac{1}{2}\kappa_{3}\Sigma=\tau_{30}\frac{\xi^{3/2}}{Y^{12}}\Biggr{[}1+\tau_{31}\frac{Y^{10}}{\xi^{2}}\Biggr{]}^{1/2},$ (38)
respectively, where
$\tau_{10}$ $=$ $\frac{1}{\sqrt{2}}\frac{\dot{m}}{\alpha_{\rm SS}}\frac{1}{(C_{\rm gap}a_{0})^{1/2}},$
$\tau_{20}$ $=$ $\frac{1}{2^{3/4}}\left(\frac{\dot{m}}{\alpha_{\rm SS}}\right)^{3/2}\frac{1}{(C_{\rm gap}a_{0})^{5/4}}\left(\frac{\kappa_{\rm ff,0}}{r_{\rm S}}\right)^{1/2}\frac{x_{e}}{\kappa_{\rm es}}\left(\frac{1+x_{e}}{T_{0}}\right)^{7/4},$
$\tau_{30}$ $=$ $\frac{1}{2}\left(\frac{\dot{m}}{\alpha_{\rm SS}}\right)^{2}\frac{1}{(C_{\rm gap}a_{0})^{2}}\frac{\kappa_{\rm ff,0}}{\kappa_{\rm es}^{2}}\frac{1}{r_{\rm S}}\left(\frac{1+x_{e}}{T_{0}}\right)^{7/2}\left[x_{e}^{2}+(1-x_{e})\frac{\kappa_{\rm bf,0}}{\kappa_{\rm ff,0}}\right],$
$\tau_{31}$ $=$ $\sqrt{2}\frac{\alpha_{\rm SS}}{\dot{m}}(C_{\rm gap}a_{0})^{3/2}\frac{\kappa_{\rm es}^{2}r_{\rm S}}{\kappa_{\rm ff,0}}\left(\frac{T_{0}}{1+x_{e}}\right)^{7/2}\left[\frac{1}{x_{e}^{2}+(1-x_{e})(\kappa_{\rm bf,0}/\kappa_{\rm ff,0})}\right].$
3.2 Energy equation
Combining equation (5) with equation (28) yields the radiative cooling rate as a function of $\xi$.
$Q_{\rm rad}=\frac{32}{3}\frac{\sigma{T_{\rm c}^{4}}}{\tau}=\frac{32}{3}\frac{1}{(1+x_{e})^{4}}\frac{\sigma{T_{\rm 0}^{4}}}{\tau}\frac{Y^{8}}{\xi^{4}}.$ (39)
Also, combining equation (8) with equation (27) yields the viscous heating rate as a function of $\xi$.
$Q_{\rm vis}=\frac{3}{4\pi}\frac{GM\dot{M}}{r_{\rm in}^{3}}\frac{1}{\xi^{3}}.$ (40)
Next, the irradiation heating rate of the CBD is given by equation (16) of Lee et al. (2024) as
$Q_{\rm irr}=\frac{A_{1}L_{1}}{2\pi r_{\rm in}^{2}}\frac{1}{\xi}\Biggr{[}(1+Q_{12})\frac{dY}{d\xi}-\frac{\beta_{1}+Q_{12}\beta_{2}}{\xi^{2}}\left(\frac{Y}{\xi}-\frac{1}{2}\frac{dY}{d\xi}\right)\Biggr{]},\$ (41)
where $A_{1}$ and $A_{2}$ are the absorption fractions of the incident radiation energy from the CPD and the CSD, respectively, $Q_{12}=A_{2}L_{2}/(A_{1}L_{1})$, and $\beta_{1}$ and $\beta_{2}$ are the parameters given by
$\beta_{1}=\frac{3}{2}\frac{1}{C_{\rm gap}^{2}}\frac{q^{2}}{(1+q)^{2}},\,\,\beta_{2}=\frac{3}{2}\frac{1}{C_{\rm gap}^{2}}\frac{1}{(1+q)^{2}},$ (42)
respectively. Note that $\beta_{1}$ and $\beta_{2}$ are less than unity for $q\leq 1$. Since $A_{1}$ and $A_{2}$ are unlikely to differ significantly for two minidisks in similar physical states, we assume that $A_{1}=A_{2}=A$, resulting in $Q_{12}=q$.
Therefore, with the dimensionless variables, the energy balance equation $Q_{\rm vis}+Q_{\rm irr}=Q_{\rm rad}$ is expressed as
$\frac{3}{4\pi}\frac{GM\dot{M}}{r_{\rm in}^{3}}\frac{1}{\xi^{3}}+\frac{AL_{1}}{2\pi r_{\rm in}^{2}}\frac{1}{\xi}\left[(1+Q_{12})\frac{dY}{d\xi}-\frac{\beta_{1}+Q_{12}\beta_{2}}{\xi^{2}}\left(\frac{Y}{\xi}-\frac{1}{2}\frac{dY}{d\xi}\right)\right]=\frac{32}{3}\frac{1}{(1+x_{e})^{4}}\frac{\sigma{T_{\rm 0}^{4}}}{\tau}\frac{Y^{8}}{\xi^{4}}.$ (43)
Model Opacity ($\kappa$) Optical depth ($\tau$) Temperature relation Outer boundary condition ($Y_{\rm out}$) Remarks
Model 0 -- $16/3$ $T_{\rm s}=T_{\rm c}$ Equation (29) of Lee et al. (2024) Lee et al. (2024)
Model 1 $\kappa_{\rm es}$ Equation (36) $T_{\rm s}\neq T_{\rm c}$ Equation (50) --
Model 2 $\sqrt{\kappa_{\rm es}\kappa_{\rm ff}}$ Equation (37) $T_{\rm s}\neq T_{\rm c}$ Equation (51) --
Model 3 $\sqrt{\kappa_{\rm a}\left(\kappa_{\rm a}+\kappa_{\rm es}\right)}$ Equation (38) $T_{\rm s}\neq T_{\rm c}$ Equation (LABEL:eq:yout3) --
Table 1: Summary of the models incorporating three different opacity prescriptions. The columns, from left to right, list the model identifier, the adopted opacity $\kappa$, the expression for the optical depth $\tau$, the relation between the surface temperature $T_{\rm s}$ and midplane temperature $T_{\rm c}$, the outer boundary condition $Y_{\rm out}$, and additional remarks. Model 0 reproduces the fiducial case from Lee et al. (2024), assuming $T_{\rm s}=T_{\rm c}$ without employing any physical opacity. Model 1 adopts pure electron scattering opacity. Model 2 describes the regime where electron scattering dominates over free-free absorption. Model 3 includes an effective opacity accounting for both free-free and bound-free absorption and electron-scattering.
3.3 Basic equation
Equation (43) leads to the differential equation to determine the radial distribution of the disk aspect ratio as
$\frac{dY}{d\xi}=\left[\frac{16}{3}\frac{\alpha}{\tau}\frac{1}{(1+x_{e})^{4}}\frac{Y^{8}}{\xi^{3}}+\beta\frac{Y}{\xi^{3}}-\gamma\frac{1}{\xi^{2}}\right]\left[1+\frac{\beta}{2}\frac{1}{\xi^{2}}\right]^{-1},$
where we introduced the following two parameters:
$\alpha$ $=\frac{1}{A}\frac{L_{0}}{L_{1}+L_{2}},$ (45)
$\beta$ $=\beta_{1}\frac{L_{1}}{L_{1}+L_{2}}+\beta_{2}\frac{L_{2}}{L_{1}+L_{2}}$ (46)
with the normalization luminosity defined by
$L_{0}$ $\equiv 4\pi r_{\rm in}^{2}\sigma{T}_{0}^{4}$
$\sim 1.3\times 10^{56}\,{\rm ergs^{-1}}\left(\frac{M}{100\,M_{\odot}}\right)^{2}\left(\frac{C_{\rm gap}}{2}\right)^{-2}\left(\frac{a_{0}}{1000}\right)^{-2}.$
Here, $\alpha$ is the ratio of the blackbody luminosity at the inner edge of the CBD to the total irradiation luminosity from the two minidisks, while $\beta$ is an average of $\beta_{1}$ and $\beta_{2}$, weighted by the ratio of the total irradiation to respective irradiation luminosities. Also, $\gamma$ is defined as
$\gamma$ $=\frac{3}{2}\frac{1}{AL_{1}}\frac{1}{1+Q_{12}}\frac{GM\dot{M}}{r_{\rm in}}=\frac{9}{2}\frac{1}{A}\frac{1}{C_{\rm gap}a_{0}}$ (48)
$\sim 2.3\times 10^{-2}\left(\frac{A}{0.1}\right)^{-1}\left(\frac{C_{\rm gap}}{2}\right)^{-1}\left(\frac{a_{0}}{1000}\right)^{-1}.$
Since equation (48) indicates that $\gamma/\xi^{2}$ goes to zero for $\xi\gg 1$, equation (44) is reduced to
$\frac{dY}{d\xi}=\left[\frac{16}{3}\frac{\alpha}{\tau}\frac{1}{(1+x_{e})^{4}}\frac{Y^{8}}{\xi^{3}}+\beta\frac{Y}{\xi^{3}}\right]\left[1+\frac{\beta}{2}\frac{1}{\xi^{2}}\right]^{-1}.$ (49)
This equation is analytically solvable for a given optical depth and a fixed value of $x_{e}$, yielding the outer boundary conditions to solve equation (44) numerically:
$Y_{\rm out}\left(\xi;\tau_{10}\right)$ $=$ $\left(\frac{\tau_{10}}{32\alpha}\right)^{1/9}\,\xi_{\rm out}^{1/6}\left[1-\frac{11}{42}\frac{\beta}{\xi_{\rm out}^{2}}\right],$ (50)
$Y_{\rm out}\left(\xi;\tau_{20}\right)$ $=$ $\left(\frac{15}{448}\frac{\tau_{20}}{\alpha}\right)^{1/14}\,\xi_{\rm out}^{5/28}\left[1-\frac{17}{84}\frac{\beta}{\xi_{\rm out}^{2}}\right],$ (51)
$Y_{\rm out}\left(\xi;\tau_{30}\right)$ $=$ $\left(\frac{21}{608}\frac{\tau_{30}}{\alpha}\right)^{1/19}\,\xi_{\rm out}^{7/38}\left[1-\frac{69}{418}\frac{\beta}{\xi_{\rm out}^{2}}\right],$
where $\xi_{\rm out}=r_{\rm out}/r_{\rm in}$ with $r_{\rm out}$ being the outer radius of the CBD. For details, see Appendix A, where approximate solutions for $\xi\gg 1$ are given by equations (A26)-(A28). The explicit form of $\tau_{10},\tau_{20},\tau_{30}$ are given by equations (A4)-(A6), respectively. The radial temperature profile is then obtained by numerically solving equation (28) with the numerical solutions for equation (44).
Our model has seven parameters: $M$, $\dot{m}$, $a_{0}$, $C_{\rm gap}$, $q$, $A$, and $\xi_{\rm out}$. Among them, the following five parameters are fixed throughout the paper as $\dot{m}=1$, $a_{0}=1000$, $C_{\rm gap}=2$, $q=1$, and $A=0.1$. The dependencies of the CBD temperature and spectra on the remaining two parameters are examined in the next section.
4 Results
In this section, we provide the numerical solutions for the radial temperature distribution of the CBD and the corresponding spectrum, and discuss how they depend on the opacity models. We also calculate the triple disk spectrum and compare it with the observational detection limits.
4.1 Circumbinary Disk Temperature Profiles
The midplane temperature, $T_{\mathrm{c}}(r)$, is obtained by numerically integrating equation (44) for the CBD aspect ratio, inserting the solution into equation (28), and solving it. The corresponding surface temperature, $T_{\mathrm{s}}(r)$, is then obtained from the radiative-equilibrium relation given by equation (7). We evaluate four opacity prescriptions (Models 0–3; see Table 1) in order to isolate the contribution of electron-scattering, free–free, and bound–free absorption to the thermal structure of CBD.
Figure 2 compares $T_{\rm c}$ and $T_{\rm s}$ for the four models. Panel (a) shows that due to the opacity effect, the midplane temperatures $T_{\mathrm{c}}$ of Models 1-3 is significantly higher than that of Model 0, except in the outermost regions for Models 1 and 2. Among Models 1-3, the difference of $T_{\mathrm{c}}$ increases with $\xi$, as both the free-free and bound-free absorptions increase with decrease of disk temperature. In contrast, panel (b) demonstrates that the surface temperatures are almost indistinguishable for $\xi\lesssim 10$, where viscous heating dominates and is effectively independent of opacity. Irradiation affects only layers near the surface; hence subtle, model-dependent differences appear at larger radii, where irradiation becomes comparable to, or exceeds, viscous heating.
Disk opacity alters the midplane temperature profile more substantially than the surface temperature profile. Since the emergent spectrum is determined by $T_{\rm s}$, only modest spectral differences between the models are expected, primarily in the outer regions where irradiation heating dominates. The subsequent section quantifies these differences and explores their observational consequences.
Figure 2: Radial temperature profiles of the CBD for the four opacity models in Table 1. Panels (a) and (b) show the midplane temperature $T_{\mathrm{c}}$ and the surface temperature $T_{\mathrm{s}}$, respectively. Both quantities are plotted against the normalized radius $\xi$ on logarithmic axes. Solid curves correspond to Model 0 (black), Model 1 (blue), Model 2 (green), and Model 3 (red).
4.2 Circumbinary Disk Spectra
Because the CBD is optically thick in the vertical direction, its surface emits local blackbody radiation with spectral intensity given by
$I_{\nu}=\frac{2h}{c^{2}}\frac{\nu^{3}}{\exp(h\nu/kT_{\rm s})-1},$
where $h$ is the Planck constant, $k$ is the Boltzmann constant, and $\nu$ is the frequency. The flux density to be emitted from the whole CBD surface is then given by (Kato et al., 2008)
[…]
where $\delta$ is the inclination angle of the disk and $D$ is the distance between the source and the earth. In the following we adopt $\delta=0$ unless otherwise noted.
Figure 3: SEDs of the CBD for the four opacity models listed in Table 1. Panels (a)–(c) correspond to total binary masses of $100\,M_{\odot}$, $10\,M_{\odot}$ (fiducial), and $1000\,M_{\odot}$, respectively. Each panel plots the normalized luminosity $\nu L_{\nu}/L_{\rm Edd}$ versus frequency on logarithmic axes. Solid curves indicate Model 0 (black), Model 1 (blue), Model 2 (green), and Model 3 (red); the dashed black curve reproduces Model 0 without external irradiation.
Figure 4: Same format as Figure 3 but for different outer disk radii, varying the outer radius $\xi_{\rm out}$ while keeping the binary mass fixed at $M=100\,M_{\odot}$. Panels (a)–(c) adopt $\xi_{\rm out}=10^{3}$, $10^{4}$, and $10^{5}$, respectively; all other parameters are identical to those in Figure 3.
Figure 3 illustrates how the emergent CBD spectrum responds to both the total binary mass and the choice of opacity prescription. For our fiducial case of $M=100\,M_{\odot}$ (Panel a), the spectrum shows the double-peaked shape, with the low-frequency peak produced by reprocessed irradiation in the outer disk and the high-frequency peak set by viscous heating in the inner regions. Introducing bound-free and free-free absorption (Model 3, red) flattens the IR-optical continuum; the height of the low-frequency peak remains within $\lesssim 0.1$ dex of the opacity-free spectrum (Model 0, black); Models 1 and 2 occupy intermediate positions and their spectral shapes are closer to that of Model 0, confirming that free-free absorption is the principal driver of this deviation.
The dashed black line represents the spectrum of the CBD without irradiation heating. In this case, the low-frequency peak is entirely absent, and the spectrum is dominated by the inner disk’s thermal emission powered purely by viscous dissipation. This remarkable difference underscores the role of irradiation in shaping the outer disk emission; without it, the surface temperature in the outer regions remains too low to contribute significantly to the IR-optical band. The contrast between the dashed and solid curves highlights that the low-frequency enhancement in the irradiated models arises from externally reprocessed radiation, rather than from local viscous heating.
Varying the total binary mass primarily shifts the entire SED along the frequency axis. For a lower-mass system ($M=10\,M_{\odot}$; Panel b), both the irradiation-induced IR peak and the viscously heated ultraviolet (UV) shoulder shift to higher frequencies. In contrast, a higher-mass system ($M=1000\,M_{\odot}$; Panel c) produces a cooler and more extended CBD, moving both features to lower frequencies.
At high frequencies ($\nu\gtrsim 10^{15}\,\mathrm{Hz}$), all four models yield nearly identical spectra. In this regime, viscous heating dominates, and the surface temperature is governed by the local energy dissipation rate, making the spectral output largely independent of the assumed opacity.
At lower frequencies, the IR peak shows a non-monotonic trend in flux across the models. Model 0, which assumes $T_{\mathrm{s}}=T_{\mathrm{c}}$ and ignores all physical opacity sources, produces the brightest spectrum. This is because the surface temperature is set equal to the higher midplane temperature, and radiation escapes without being absorbed. Among the models that include opacity, Model 3 gives the strongest IR flux, followed by Model 2 and then Model 1. This order reflects differences in surface temperature profiles (see Figure 2), which are caused by how efficiently each model absorbs and thermalizes the incoming radiation in the outer part of the disk.
In Model 3, the inclusion of bound–free absorption greatly increases the Rosseland mean opacity in the temperature range $T\sim 10^{3}$$10^{4}\,\mathrm{K}$, which coincides with the irradiated outer disk regions where the disk flares. This enhances the local trapping of irradiating photons and raises the surface temperature, leading to stronger IR re-emission than in Model 2, where only free–free opacity is present. In contrast, electron scattering only in Model 1 does not effectively trap irradiating photons due to the weaker disk flaring, so the surface remains cooler and the IR flux is lower.
These trends indicate that observational constraints on CBD opacity sources are most effective in the IR for high-mass binaries and in the optical for low-mass binaries, where the spectral differences among models are most pronounced.
In Panel (b) of Figure 3, the spectrum of Model 3 (red curve) clearly exceeds those of the other three models across the mid-IR through the optical and near-ultraviolet frequencies. Only in the far-IR (frequencies $\nu\lesssim 10^{13.5}\,\mathrm{Hz}$) does Model 0 (black curve), which includes no physical absorption opacity, slightly surpass Model 3. Thus, within a single panel, a two-tier behavior emerges: Model 0 leads only at the lowest frequencies, while Model 3 dominates elsewhere. Furthermore, when the SEDs of Model 3 are compared across the three mass cases (Panels a, b, and c), the $10\,M_{\odot}$ case exhibits the highest normalized luminosity.
This behavior arises from the combination of the CBD’s temperature profile and the reprocessing of irradiation from the two minidisks surrounding the binary. In the $10\,M_{\odot}$ system, the outer disk cools to temperatures below $10^{4}\,$, where hydrogen is partially ionized and bound–free opacity becomes especially strong. High-energy photons emitted by the two minidisks are absorbed in these outer layers, thermalized, and then re-emitted at IR wavelengths, producing a pronounced IR peak. Meanwhile, the inner disk remains hot enough that its surface temperature does not drop significantly, enhancing the mid- and high-frequency emission.
The apparent “triple peaked” shape in the SED of Model 3 is in fact a modification of the basic double-peaked SED structure. The two primary peaks, one from viscous heating in the inner CBD and the other from reprocessed emission from the irradiated outer disk, are supplemented by a smaller subpeak in the mid-frequency range. This subpeak arises because bound–free absorption in Model 3 distorts the irradiated component of the spectrum, creating a visible peak between the main IR peak and the UV shoulder. Therefore, it is most appropriate to describe the SED as fundamentally double-peaked, with an additional mid-frequency peak caused by bound–free opacity effects.
Figure 4 shows how the CBD spectrum changes when the outer disk radius $r_{\rm out}$ varies. Panels (a), (b), and (c) correspond to $r_{\rm out}=10^{4}\,r_{\rm in}$ (fiducial case), $10^{3}\,r_{\rm in}$, and $10^{5}\,r_{\rm in}$, respectively. The black, blue, green, and red curves represent Models 0, 1, 2, and 3, respectively. As $r_{\rm out}$ increases, the IR peak becomes stronger. Panel (c), which adopts the largest outer radius, shows the most prominent IR peak. This is because a larger outer radius allows the two minidisks to irradiate a wider region of the outer CBD. The radiation from the minidisks is absorbed and then re-emitted as IR light, increasing the total IR luminosity. In contrast, Panel (b), for the smallest outer radius, shows a much weaker IR peak due to its limited irradiated area.
The spectral differences between the models also become more noticeable in the IR region as $r_{\rm out}$ increases. This is because the opacities adopted in each model mainly affect the efficiency of reprocessing the irradiating photons. When irradiation covers a larger portion of the disk, these opacity effects become more significant and lead to clearer differences among the models. Therefore, both the size of the CBD and the strength of irradiation are key to understanding how different types of opacity—such as electron scattering, free–free absorption, and bound–free absorption—influence the emergent spectrum.
4.3 Triple disk spectra
We compute the SED of a triple disk system consisting of the CPD, the CSD, and the surrounding CBD. The spectral luminosity of each minidisk is calculated using the relation $L_{\nu,i}=4\pi D^{2}S_{\nu,i}$, where $S_{\nu,i}$ is the flux density emitted from the CPD ($i=1$) or the CSD ($i=2$). This flux density is computed using the same integrand as in equation (53), but it is integrated over a distinct radial range, from the ISCO radius, $r_{{\rm in},i}=6GM_{i}/c^{2}$, to the outer edge of each disk. The outer edge is defined as a fixed fraction (typically $\sim 0.5$) of the Hill radius, which represents the tidal truncation limit for each minidisk. This choice reflects the results of numerical simulations showing that tidal truncation limits the disk extent well within the Hill sphere (e.g., Martin & Lubow 2011; Paschalidis et al. 2021). The temperature profile of each minidisk follows the standard thin-disk scaling, $T_{i}(r)\propto r^{-3/4}$, where $r_{{\rm in},i}\leq{r}\leq{r_{{\rm out},i}}$.
Figure 5: Comparison of the combined SEDs of the triple disk, consisting of the CPD, the CSD, and the CBD surrounding them. The vertical and horizontal axes represent the normalized luminosity and frequency, respectively, both plotted on a logarithmic scale. For all models, the same CPD and CSD parameters are used, and differences in the triple disk spectra arise solely from the CBD component. The solid black, blue, green, and red lines denote the total SEDs for Models 0, 1, 2, and 3, respectively. The dashed black line shows the triple disk spectrum without irradiation heating, while the dotted black line shows the combined SED from the CPD and CSD only, excluding the CBD contribution. The other CBD parameters are identical to those adopted in panel (a) of Figure 3 for all models.
Figure 5 illustrates how the combined SED of the triple disk system responds to different opacity prescriptions applied to the CBD. Since the CPD and CSD parameters are held fixed across all models, the differences among the spectra reflect only the variation in the CBD treatment. In all cases, the triple disk SED exhibits a double-peaked structure: a high-frequency (X-ray) peak originating from the hot inner regions of the CPD and CSD, and a lower-frequency (IR) peak caused by reprocessed irradiation in the outer CBD, as shown by Lee et al. (2024). Note that the IR component shows significant model dependence due to the opacity-dependent surface temperature profile of the CBD.
The dashed black curve in Figure 5 represents the total SED when the irradiation heating term is omitted from the CBD, thereby isolating the role of viscous heating alone. Compared to the irradiated cases, the IR peak is significantly suppressed, confirming that irradiation is essential for generating the low-frequency excess. The dotted line shows the spectrum produced solely by the CPD and CSD without any CBD contribution. Notably, this curve produces the X-ray peak but lacks the IR component entirely. This comparison highlights the distinct spectral signatures arising from each disk component: the CPD and CSD dominate the high-energy emission, while the CBD is the sole contributor to the IR peak. The deviation between models becomes more prominent in the IR regime, demonstrating the impact of different opacity prescriptions on the CBD’s reprocessing efficiency.
4.4 Observational implications
To evaluate the detectability of the model spectra, we compare the predicted flux densities with the sensitivity limits of major current telescopes. These include Subaru/HSC and MOIRCS in the optical and near-IR (NIR) bands, JWST/NIRCam in the mid-to-NIR, and Swift/XRT in the X-ray band. Figure 6 shows the triple disk spectra for Models 0 and 3 at source distances of 10 Mpc and 100 Mpc, overlaid with the flux detection thresholds of these instruments. All other model parameters are fixed at their fiducial values ($M=100\,M_{\odot}$, $r_{\rm out}=10^{4}\,r_{\rm in}$).
At $D=100\,{\rm Mpc}$, even Model 3, which is our most optically thick model, falls below the detection thresholds in the optical and NIR bands. The only potentially observable signal in this case is the X-ray emission, originating from the two minidisks (CPD and CSD), which is detectable with Swift/XRT. In contrast, at $D=10\,{\rm Mpc}$, Model 3 exhibits a distinct low-frequency excess due to the irradiated CBD, making it detectable with JWST and also marginally detectable with the Subaru/HSC $r2$ filter. This contrast highlights the role of distance in the detectability of the CBD contribution in the IR-optical regime.
The fact that irradiation-induced CBD features emerge in the optical/NIR only at nearby distances ($\lesssim 10\,{\rm Mpc}$) sets practical constraints on follow-up campaigns. In particular, detecting the IR peak requires deep imaging of relatively nearby sources with multiple filters to disentangle the spectral shape. These features may appear as broad, red excesses in the SED, distinguishable from the steeper blue continuum of minidisk-dominated emission.
Based on this, we propose a targeted follow-up strategy: Upon detection of a bright X-ray flare by Swift/XRT or other X-ray facilities, especially if a distance estimate places the source within $\lesssim 100$ Mpc, NIR imaging should be promptly initiated using telescopes such as Subaru or JWST. If a counterpart is found, a detailed spectro-photometric observation across multiple bands can constrain the SED slope and test for IR flattening predicted by our Model 3. Such multi-band follow-up would be key to identifying reprocessed emission from irradiated CBDs.
Figure 6: Comparison of Subaru, JWST, and Swift/XRT sensitivity limits with theoretical triple disk spectra at different source distances. The vertical axis shows the flux density (in erg s-1 cm-2 Hz-1), and the horizontal axis represents the frequency (in Hz); both are plotted on logarithmic scales. All models adopt fiducial parameters of $M=100\,M_{\odot}$ and $r_{\rm out}=10^{4}\,r_{\rm in}$. The solid black line denotes Model 0, while the solid red and green lines represent Model 3 at distances of $D=10$ Mpc and $D=100$ Mpc, respectively. The dashed horizontal curves indicate the flux sensitivity limits of JWST/NIRCam filters (F [^4] W through F070W), Subaru/MOIRCS and HSC filters ($Y$, $z$, $i2$, $r2$, and $g$), and Swift/XRT in the 0.2–10 keV X-ray band. This comparison highlights the observability of CBD features in the IR-optical-X-ray bands and demonstrates that, for $D=100$ Mpc, only the X-ray emission (primarily from the CPD and CSD) is detectable, whereas at $D=10$ Mpc, the irradiated CBD in Model 3 becomes marginally detectable in the Subaru/HSC $r2$ band.
Figure 7: Rosseland mean opacities at a fixed density of $\rho=10^{-10}\,\mathrm{g\,cm^{-3}}$. The dashed green, magenta, and orange lines represent the contributions from free-free absorption, bound-free absorption, and electron scattering for hydrogen gas, respectively. The solid black line represents the total analytical opacity, calculated as $\kappa_{\rm ff}+\kappa_{\rm bf}+x_{e}\,\kappa_{\rm es}$, where the ionization fraction $x_{e}$ is determined self-consistently from the Saha equation. The electron scattering term $\kappa_{\rm es}$ is multiplied by $x_{e}$ to account for its dependence on free electrons. The solid blue line indicates the tabulated opacity from the Opacity Project (Seaton, 2005), an international collaboration to calculate the extensive atomic data required to estimate stellar envelope opacities, and from (Semenov et al., 2003), including dust opacity based on normal silicate grains under the assumption of homogeneous spheres.
5 Discussion
This section discusses two key caveats of our models. First, we examine how realistic opacity prescriptions affect the emergent spectrum from an irradiated CBD, focusing on discrepancies between our models and more detailed opacities available in tabulated forms. Second, we assess the detectability of GWs emitted from such systems by current and future GW observatories, based on the binary orbital parameters.
In Figure 7, we compare the Rosseland mean opacity computed from our models—including free-free absorption, bound-free absorption, and electron scattering processes—with the tabulated values from the Opacity Project (Seaton, 2005), which provides extensive atomic data for stellar envelopes, and from Semenov et al. (2003), widely used in the context of protoplanetary disk modeling. Our model employs cross-section formulae and ionization fractions derived from the Saha equation, assuming pure hydrogen gas at a fixed density of $\rho=10^{-10}~{}{\rm g\,cm^{-3}}$. The comparison reveals significant discrepancies, particularly in the temperature range of $T\sim 10^{3}$$10^{4}$ K, where the tabulated Rosseland opacities are several orders of magnitude lower than those predicted by our analytical prescription that includes the bound-free absorption. This divergence arises because the tabulated opacities incorporate more realistic astrophysical conditions, such as grain sublimation, grain growth, and composition-dependent absorption features, all of which are absent from our simplified hydrogen-only framework.
These differences in opacity have direct consequences for the emergent spectrum from irradiated CBDs. In the temperature range of $T\sim 10^{3}$$10^{4}$ K, the reduced opacity in the tabulated models is expected to enhance radiative cooling, allowing energy to be transported more efficiently to the disk surface and resulting in a higher surface temperature. Consequently, the overall spectral intensity, particularly in the far-IR and sub-millimeter bands dominated by reprocessed irradiation, increases.
As a future direction, it will be essential to incorporate opacity tables (e.g., Semenov et al. 2003; Seaton 2005) directly into irradiated disk models, enabling systematic computation and comparison of CBD spectra under more realistic astrophysical conditions. This will provide a more robust framework for interpreting multi-wavelength observations and constraining the physical parameters of circumbinary environments.
If the binary system is undergoing orbital decay due to GW emission, the emitted GW signal is characterized by its frequency and strain amplitude, both determined by the binary’s orbital parameters. In our fiducial model, the GW frequency and strain are given by (Moore et al., 2015)
$f_{\rm GW}$ $=$ $\frac{\Omega}{\pi}\sim 7.2\times 10^{-3}\,{\rm Hz}\left(\frac{M}{100\,M_{\odot}}\right)^{-1},$ (54)
$h_{\rm c}$ $=$ $\frac{2(G\mathcal{M})^{5/3}(\pi f_{\rm GW})^{2/3}}{c^{4}D}\sim 1.2\times 10^{-22}\left(\frac{M}{100\,M_{\odot}}\right)\left(\frac{D}{10\,{\rm Mpc}}\right)^{-1},$ (55)
respectively, where $\mathcal{M}=(M_{1}M_{2})^{3/5}/(M_{1}+M_{2})^{1/5}$ is the chirp mass, and $D$ is the distance to the source.
[…]
Therefore, GWs emitted from the BBH systems modeled in this work are unlikely to be detected by current or near-future GW observatories. In future work, we aim to extend our model to explore systems with either higher frequency or stronger GW emission, such as stellar-mass BBHs detectable by LIGO/Virgo/KAGRA or supermassive BBH systems within LISA’s sensitivity range. Such extensions will allow us to construct and analyze physically consistent models that are potentially observable through both GW and EM channels.
6 Conclusions
We have investigated the effect of opacity prescriptions on the thermal and spectral properties of irradiated CBDs surrounding BBH systems. Our models incorporated temperature- and density-dependent hydrogen opacities, specifically considering electron scattering, free-free absorption, and bound-free absorption. By numerically solving the local energy balance equation with self-consistent outer boundary conditions, we analyzed how different opacity prescriptions shape the temperature profiles and resultant spectra of the CBD. Our main findings are summarized as follows:
  1. 1.
    The CBD surface temperature profiles remain largely consistent across different opacity models in the regions dominated by viscous heating. However, significant difference occurs in the midplane temperature profiles due to the substantial influence of disk opacity.
  2. 2.
    Opacity effects predominantly manifest in the lower-frequency regime of the CBD spectrum, where external irradiation from minidisks dominates heating. These effects are most pronounced in the IR and optical bands.
  3. 3.
    Inclusion of analytical hydrogen absorption opacities notably flattens the continuum emission in the IR-optical range, creating an additional spectral hump distinct from the standard double-peaked profile identified in the earlier simpler models (Lee et al., 2024).
  4. 4.
    Despite the difference in the opacity prescriptions, the combined spectrum of the triple disk system, which consists of the CBD and two minidisks, retains a characteristic double-peaked structure. The high-frequency X-ray peak primarily arises from the inner regions of the minidisks, while the low-frequency IR peak results from irradiation reprocessing in the CBD.
  5. 5.
    Observational feasibility analysis demonstrates that current instruments like Subaru/HSC, JWST/NIRCam, and Swift/XRT provide adequate sensitivity for distinguishing among our opacity models, particularly at distances less than $\sim 10$ Mpc. Hence, coordinated multi-band follow-up observations following X-ray flares could effectively test our model predictions and discern the physical conditions in irradiated CBDs.
Our results underscore the critical importance of accurately modeling opacity in astrophysical disk systems, highlighting the necessity for incorporating more realistic opacity prescriptions in future theoretical studies. This advancement will significantly enhance our ability to interpret electromagnetic signals associated with gravitational wave events from binary black hole mergers.
Acknowledgments
S.B. acknowledges support from the Basic Science Research Program through the National Research Foundation of Korea (NRF), funded by the Ministry of Education (grant RS-2024-00460704). K.H. and S.B. were also supported by the same program (grant 2020R1A2C1007219). This work was conducted during the research year of Chungbuk National University in 2021. This work was partially supported by NSF grant PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP), and by the Grant-in-Aid for Scientific Research from MEXT/JSPS of Japan (grant JP21K03619 to A.T.O.).
Appendix A Analytical solutions for $\beta\neq 0$
A.1 Optical depth prescriptions
In this section, we describe the optical depth prescriptions. Equations (36)-(38), which show the optical depths as a function of $Y$ and $\xi$, are rewritten as
$\tau_{1}$ $=$ $\frac{1}{2}\kappa_{\rm es}\Sigma=\tau_{10}\frac{1}{Y^{2}\xi^{1/2}},$ (A1)
$\tau_{2}$ $=$ $\frac{1}{2}\sqrt{\kappa_{\rm es}\kappa_{\rm ff}}\;\Sigma=\tau_{20}\,\frac{\xi^{1/2}}{Y^{7}},$ (A2)
$\tau_{3}$ $=$ $\frac{1}{2}\sqrt{\kappa_{\rm a}\!\left(\kappa_{\rm a}+\kappa_{\rm es}\right)}\;\Sigma=\frac{1}{2}\kappa_{3}\Sigma=\tau_{30}\frac{\xi^{3/2}}{Y^{12}}\,\Biggr{[}1+\tau_{31}\frac{Y^{10}}{\xi^{2}}\Biggr{]}^{1/2}\approx\tau_{30}\frac{\xi^{3/2}}{Y^{12}}+\frac{\tau_{10}}{2}\frac{1}{\xi^{1/2}Y^{2}}\approx\tau_{30}\frac{\xi^{3/2}}{Y^{12}},$ (A3)
where $x_{e}=1$ for $\tau_{20}$ and $x_{e}=0$ for $\tau_{31}$ and $\tau_{32}$ are adopted so that
$\tau_{10}$ $=$ $\frac{1}{\sqrt{2}}\frac{\dot{m}}{\alpha_{\rm SS}}\frac{1}{(C_{\rm gap}a_{0})^{1/2}},$ (A4)
$\tau_{20}$ $=$ $\frac{1}{2^{3/4}}\left(\frac{\dot{m}}{\alpha_{\rm SS}}\right)^{3/2}\frac{1}{(C_{\rm gap}a_{0})^{5/4}}\left(\frac{\kappa_{\rm ff,0}}{r_{\rm S}}\right)^{1/2}\frac{x_{e}}{\kappa_{\rm es}}\left(\frac{1+x_{e}}{T_{0}}\right)^{7/4}=2\left(\frac{\dot{m}}{\alpha_{\rm SS}}\right)^{3/2}\frac{1}{(C_{\rm gap}a_{0})^{5/4}}\left[\left(\frac{\kappa_{\rm ff,0}}{\kappa_{\rm es}^{2}r_{\rm S}}\right)T_{0}^{-7/2}\right]^{1/2},$ (A5)
$\tau_{30}$ $=$ $\frac{1}{2}\left(\frac{\dot{m}}{\alpha_{\rm SS}}\right)^{2}\frac{1}{(C_{\rm gap}a_{0})^{2}}\frac{\kappa_{\rm ff,0}}{\kappa_{\rm es}^{2}}\frac{1}{r_{\rm S}}\left(\frac{1+x_{e}}{T_{0}}\right)^{7/2}\left[x_{e}^{2}+(1-x_{e})\frac{\kappa_{\rm bf,0}}{\kappa_{\rm ff,0}}\right]=\frac{1}{2}\left(\frac{\dot{m}}{\alpha_{\rm SS}}\right)^{2}\frac{1}{(C_{\rm gap}a_{0})^{2}}\frac{\kappa_{\rm bf,0}}{\kappa_{\rm es}^{2}r_{\rm S}}{T_{0}}^{-7/2},$ (A6)
$\tau_{31}$ $=$ $\sqrt{2}\frac{\alpha_{\rm SS}}{\dot{m}}(C_{\rm gap}a_{0})^{3/2}\frac{\kappa_{\rm es}^{2}r_{\rm S}}{\kappa_{\rm ff,0}}\left(\frac{T_{0}}{1+x_{e}}\right)^{7/2}\left[\frac{1}{x_{e}^{2}+(1-x_{e})(\kappa_{\rm bf,0}/\kappa_{\rm ff,0})}\right]=\sqrt{2}\frac{\alpha_{\rm SS}}{\dot{m}}(C_{\rm gap}a_{0})^{3/2}\frac{\kappa_{\rm es}^{2}r_{\rm S}}{\kappa_{\rm bf,0}}T_{0}^{7/2}.$
A.2 Analytical solutions
For each model, equation (49) is expressed as
$\left[\xi^{3}+\frac{\beta}{2}\xi\right]\frac{dY}{d\xi}$ $=$ $\left[\frac{1}{3}\frac{\alpha}{\tau_{10}}Y^{10}\xi^{1/2}+\beta\,Y\right],$ (A7)
$\left[\xi^{3}+\frac{\beta}{2}\xi\right]\frac{dY}{d\xi}$ $=$ $\left[\frac{1}{3}\frac{\alpha}{\tau_{20}}\frac{Y^{15}}{\xi^{1/2}}+\beta\,Y\right],$ (A8)
$\left[\xi^{3}+\frac{\beta}{2}\xi\right]\frac{dY}{d\xi}$ $=$ $\left[\frac{16}{3}\frac{\alpha}{\tau_{30}}\frac{Y^{20}}{\xi^{3/2}}+\beta\,Y\right],$ (A9)
where $x_{e}=1$ is adopted for Models 1 and 2, while $x_{e}=0$ is adopted for Model 3. Now we solve these differential equations by the perturbation method with $\beta$, which is much smaller than $1$. Considering that the terms with $\beta$ in the above equations are much smaller than those without $\beta$, we can expand $Y$ for $\beta\ll 1$ as
$Y=Y_{0}(\xi)+\beta{Y_{1}(\xi)}+\beta^{2}{Y_{2}(\xi)}+\cdots.$ (A10)
Substituting equation (A10) into equations (A7)-(A9), we obtain the following equations, which are approximated at the first-order of $\beta$, as
$\xi^{3}\frac{dY_{0}}{d\xi}+\beta\left(\frac{\xi}{2}\frac{dY_{0}}{d\xi}+\xi^{3}\frac{dY_{1}}{d\xi}\right)$ $=$ $\frac{1}{3}\frac{\alpha}{\tau_{10}}Y_{0}^{10}\xi^{1/2}+\beta\left(\frac{10}{3}\frac{\alpha}{\tau_{10}}Y_{0}^{9}Y_{1}\xi^{1/2}+Y_{0}\right),$ (A11)
$\xi^{3}\frac{dY_{0}}{d\xi}+\beta\left(\frac{\xi}{2}\frac{dY_{0}}{d\xi}+\xi^{3}\frac{dY_{1}}{d\xi}\right)$ $=$ $\frac{1}{3}\frac{\alpha}{\tau_{20}}Y_{0}^{15}\xi^{-1/2}+\beta\left(5\frac{\alpha}{\tau_{20}}Y_{0}^{14}Y_{1}\xi^{-1/2}+Y_{0}\right),$ (A12)
$\xi^{3}\frac{dY_{0}}{d\xi}+\beta\left(\frac{\xi}{2}\frac{dY_{0}}{d\xi}+\xi^{3}\frac{dY_{1}}{d\xi}\right)$ $=$ $\frac{16}{3}\frac{\alpha}{\tau_{30}}Y_{0}^{20}\xi^{-3/2}+\beta\left(\frac{320}{3}\frac{\alpha}{\tau_{30}}Y_{0}^{19}Y_{1}\xi^{-3/2}+Y_{0}\right).$ (A13)
For the terms with the 0th order of $\beta$, we obtain the following identities:
$\frac{dY_{0}}{d\xi}$ $=$ $\frac{1}{3}\frac{\alpha}{\tau_{10}}Y_{0}^{10}\xi^{-5/2},$ (A14)
$\frac{dY_{0}}{d\xi}$ $=$ $\frac{1}{3}\frac{\alpha}{\tau_{20}}Y_{0}^{15}\xi^{-7/2},$ (A15)
$\frac{dY_{0}}{d\xi}$ $=$ $\frac{16}{3}\frac{\alpha}{\tau_{30}}Y_{0}^{20}\xi^{-9/2},$ (A16)
yielding the power-law solutions:
$Y_{0}$ $=$ $\left(\frac{\tau_{10}}{2\alpha}\right)^{1/9}\xi^{1/6},$ (A17)
$Y_{0}$ $=$ $\left(\frac{15}{28}\frac{\tau_{20}}{\alpha}\right)^{1/14}\xi^{5/28},$ (A18)
$Y_{0}$ $=$ $\left(\frac{21}{608}\frac{\tau_{30}}{\alpha}\right)^{1/19}\xi^{7/38}.$ (A19)
For the terms with the $1$st order of $\beta$, we get the following identities:
$\frac{\xi}{2}\frac{dY_{0}}{d\xi}+\xi^{3}\frac{dY_{1}}{d\xi}$ $=$ $\frac{10}{3}\frac{\alpha}{\tau_{10}}Y_{0}^{9}Y_{1}\xi^{1/2}+Y_{0},$ (A20)
$\frac{\xi}{2}\frac{dY_{0}}{d\xi}+\xi^{3}\frac{dY_{1}}{d\xi}$ $=$ $5\frac{\alpha}{\tau_{20}}Y_{0}^{14}Y_{1}\xi^{-1/2}+Y_{0},$ (A21)
$\frac{\xi}{2}\frac{dY_{0}}{d\xi}+\xi^{3}\frac{dY_{1}}{d\xi}$ $=$ $\frac{320}{3}\frac{\alpha}{\tau_{30}}Y_{0}^{19}Y_{1}\xi^{-3/2}+Y_{0}.$ (A22)
Substituting equations (A14)-(A19) into equations (A20)-(A22) yields the following solutions:
$Y_{1}$ $=$ $-\frac{11}{42}\left(\frac{\tau_{10}}{2\alpha}\right)^{1/9}\xi^{-11/6},$ (A23)
$Y_{1}$ $=$ $-\frac{17}{84}\left(\frac{15}{28}\frac{\tau_{20}}{\alpha}\right)^{1/14}\xi^{-51/28},$ (A24)
$Y_{1}$ $=$ $-\frac{69}{418}\left(\frac{21}{608}\frac{\tau_{30}}{\alpha}\right)^{1/19}\xi^{-69/38}.$ (A25)
From equations (A17) and (A25), we obtain the approximate solutions as
$Y(\xi;\tau_{10})$ $=$ $Y_{0}+\beta\,Y_{1}=\left(\frac{\tau_{10}}{2\alpha}\right)^{1/9}\xi^{1/6}\left[1-\frac{11}{42}\frac{\beta}{\xi^{2}}\right],$ (A26)
$Y(\xi;\tau_{20})$ $=$ $Y_{0}+\beta\,Y_{1}=\left(\frac{15}{28}\frac{\tau_{20}}{\alpha}\right)^{1/14}\xi^{5/28}\left[1-\frac{17}{84}\frac{\beta}{\xi^{2}}\right],$ (A27)
$Y(\xi;\tau_{30})$ $=$ $Y_{0}+\beta\,Y_{1}=\left(\frac{21}{608}\frac{\tau_{30}}{\alpha}\right)^{1/19}\xi^{7/38}\left[1-\frac{69}{418}\frac{\beta}{\xi^{2}}\right].$ (A28)
with fixed values of $\alpha$ and $\beta$.